### v3

# Collaborative Face Experts Fusion in Video Generation: Boosting Identity Consistency Across Large Face Poses

Abstract
Current video generation models struggle with identity preservation under large face poses, primarily facing two challenges: the difficulty in exploring an effective mechanism to integrate identity features into DiT architectures, and the lack of targeted coverage of large face poses in existing open-source video datasets. To address these, we present two key innovations. First, we propose Collaborative Face Experts Fusion (CoFE), which dynamically fuses complementary signals from three specialized experts within the DiT backbone: an identity expert that captures cross-pose invariant features, a semantic expert that encodes high-level visual context, and a detail expert that preserves pixel-level attributes such as skin texture and color gradients. Second, we introduce a data curation pipeline comprising three key components: Face Constraints to ensure diverse large-pose coverage, Identity Consistency to maintain stable identity across frames, and Speech Disambiguation to align textual captions with actual speaking behavior. This pipeline yields LaFID-180K, a large-scale dataset of pose-annotated video clips designed for identity-preserving video generation. Experimental results on several benchmarks demonstrate that our approach significantly outperforms state-of-the-art methods in face similarity, FID, and CLIP semantic alignment. Project page.
1 Introduction
Video generation has emerged as a critical technology for applications ranging from virtual avatars and interactive storytelling to realistic video editing. A critical yet unresolved challenge is maintaining the identity consistency of human characters, as human visual perception is inherently more sensitive to identity inconsistencies in human portraits than in other subjects, particularly when face poses undergo large variations. Such consistency is essential for narrative coherence in storybook generation, visual realism, and reliable character representation in multi-scene animations, as even subtle identity drift can severely undermine user immersion.
Figure 1: Top: Identity preservation under large face poses is challenging in current methods, while our approach shows clear gains in face similarity metrics. Bottom: With dynamic multi-feature injection, our proposed method achieves the best or competitive performance across all metrics.
Recent Diffusion Transformer (DiT)-based image-to-video models such as CogVideoX [yang2024cogvideox] and Wan Model [wan2025wan] achieve content consistency by injecting first-frame VAE features into the DiT backbone. To better preserve character identity, human-centric approaches like ConsisID [yuan2025consisid] further inject a frequency-aware facial representation into DiT. However, both paradigms share critical limitations: they rely on a single face feature extracted from the reference frame and inject the same static face feature uniformly across all timesteps and DiT layers, thereby limiting the model’s ability to capture discriminative facial details. This leads to unstable identity preservation under large face poses, often resulting in distorted character appearances. Furthermore, existing human-centric datasets [wang2024humanvid, li2025openhumanvid] are ill-suited for learning pose-robust identity preservation. They often feature faces of inconsistent quality, lack sufficient dynamic large-angle examples, and seldom ensure cross-frame identity consistency or sufficient facial visibility, hindering the learning of pose-invariant identity representations.
To address these gaps, we propose Collaborative Face Experts Fusion (CoFE), a hierarchical feature integration framework that overcomes the limitations of static, single-feature identity injection. Since fixed features from the first frame often fail to provide reliable identity cues under large face poses, CoFE instead dynamically fuses complementary signals from three specialized experts via a learnable gating mechanism: i) Identity Expert extracts pose-robust identity features for cross-view consistency; ii) Semantic Expert captures high-level visual concepts (e.g., hairstyle, expression); and iii) Detail Expert preserves pixel-level fidelity (e.g.,, skin texture, color gradients). By aligning expert contributions with the functional role of each DiT layer, CoFE enables robust feature learning under large face poses.
Complementing our architectural innovation, we introduce LaFID-180K (Large Face Pose for Identity Preservation Dataset), a curated large-scale collection of video clips annotated with 3D face poses. This dataset specifically targets the scarcity of high-quality, identity-consistent videos under large face poses. LaFID-180K is built through a three-stage pipeline that directly addresses key data limitations: Face Constraints ensure clear facial visibility and diverse large-pose coverage; Identity Consistency enforces temporal identity coherence via feature clustering; and Speech Disambiguation aligns video captions with lip movements by verifying speaking status, ensuring accurate annotations under pose-induced ambiguity. We evaluate our method on two distinct benchmarks: the existing VIP-Test [vip-200k] for general full-body motion, and our newly introduced LaFID-Bench, which addresses the underexplored setting of large face poses.
Our contribution can be summarized as follows:
  • We propose Collaborative Face Experts Fusion (CoFE), a dynamic, multi-source identity injection module that adaptively fuses three specialized experts across DiT layers for robust identity preservation under large face poses.
  • We introduce LaFID-180K, a dataset of 180K video clips with 3D face pose annotations, curated specifically for large-pose identity preservation via Face Constraints, Identity Consistency, and Speech Disambiguation to ensure diverse large face poses, stable identity across frames, and accurate alignment between lip movements and captions.
  • We establish LaFID-Bench for large face pose evaluation and show that CoFE achieves state-of-the-art results on both LaFID-Bench and VIP-Test, with strong gains under large face poses.
2 Related Work
Figure 2: Overview of our pipeline. Our CoFE module comprises globally shared projection layers to align different face features into the feature space of DiT and per-DiT-block expert fusion layers for dynamic feature fusion. Specifically, we employ ArcFace, CLIP, and DINOv2 as our identity, semantic, and detail encoders, respectively.
2.1 Identity-Preserving Text-to-Video Generation
In the identity-preserving text-to-video (IPT2V) task, early methods used time-consuming fine-tuning for facial identity preservation [chefer2024still, wu2024motionbooth, wei2024dreamvideo, zhong2025concat, lei2024comprehensive, ma2025controllable], driving tuning-free approaches to dominance. Pioneering works like ID-Animator [he2024idanimator] relied on static image-model features, harming dynamic identity consistency. ConsisID [yuan2025consisid] injected facial frequency signals but suffered from 2D pose mismatches; FantasyID [zhang2025fantasyid] added 3D priors but lacked adaptive feature selection. Recent methods like VACE [jiang2025vace] and SkyReels-A2 [fei2025skyreels] used reference branches but failed to dynamically prioritize cues during large motions. These show I2V needs adaptive fusion of face features aligned with motion, especially large face motions. Thus, we introduce Collaborative Face Experts Fusion (CoFE), a DiT layer-adaptive gating mechanism that dynamically fuses multiple features to overcome static fusion and enhance large-pose robustness.
2.2 Human-Centric Image-to-Video Generation
Image-to-video (I2V) generation [guo2024i2v, bharadhwaj2024gen2act, xu2025freevis] struggles to control content reliably with text alone, so recent methods use first-frame conditioning for better controllability. AnimateAnything [dai2023animateanything] and PixelDance [zeng2024make] modify U-Nets by concatenating first-frame latent features with input noise to anchor spatial features for temporal consistency. Dynamicrafter [xing2024dynamicrafter] and Moonshot [zhang2024moonshot] enhance conditioning with image cross-attention layers, injecting stronger visual cues into diffusion. Recent methods like Wan2.1, Wan2.2 [wan2025wan], and VACE [jiang2025vace] support I2V with strong performance but struggle to preserve identity across large face poses. Moreover, IPRO [shen2025identity] introduces reinforcement learning to enhance training for facial details. Generally, identity robustness under large face pose variations remains largely unaddressed in current I2V frameworks.
2.3 Human-Centric Video Datasets
Large-scale human-centric video datasets [sadoughi2015msp, chen2024panda, bain2021frozen] are foundational for human video generation, yet most prioritize body motion or general scenes over facial identity under large poses. For example, TikTok [jafarian2021learning] focuses on social dance; UCF101 [soomro2012ucf101] and NTU RGB+D [shahroudy2016ntu] on action recognition; HumanVid [wang2024humanvid] on synthetic-real blends; and OpenHumanVid [li2025openhumanvid] provides high-resolution videos but lacks large face pose diversity, facial quality filtering, and fine-grained 3D face pose annotations, limiting its suitability for pose-robust identity learning. To address it, we introduce LaFID-180K, a dataset constructed by combining OpenHumanVid [li2025openhumanvid] with carefully selected in-house videos, and curated through a unified pipeline to emphasize diverse large face poses with 3D face pose annotations.
3 Methodology: CoFE
Current image-to-video models [yang2024cogvideox, hacohen2024ltx, wan2025wan, jiang2025vace] inject identity using a single-source feature extracted from the first frame, which is then applied statically across all timesteps and network layers, typically via uniform cross-attention. This design suffers from two fundamental limitations. First, the single feature conflates identity, semantics, and details into one representation, lacking specialization. Second, its static injection cannot adapt to pose-induced appearance changes or layer-specific attention demands, leading to severe identity drift under large pose variations. To overcome this, we propose Collaborative Face Experts Fusion (CoFE), a layer-adaptive fusion module that dynamically integrates three specialized facial representations into cross-attention: pose-robust identity features, semantic attributes (e.g., hairstyle, expression), and pixel-level details (e.g., color gradients). An overview of our pipeline is shown in Fig. 2.
Figure 3: Cross-attention visualizations of different experts. Warmer colors denote higher attention strength. CoFE integrates complementary attention from Identity Expert (ArcFace), Semantic Expert (CLIP), and Detail Expert (DINOv2) to jointly enhance facial detail fidelity and identity consistency.
Motivation from Cross-Attention Patterns. To understand how different facial representations influence identity preservation, we analyze cross-attention patterns in the generator. As shown in Figure 3, each representation induces a distinct attention pattern: the Identity Expert (ArcFace [deng2019arcface]) focuses on pose-invariant landmarks such as the eyes, nose, and mouth to capture structural identity; the Semantic Expert (CLIP [radford2021learning]) attends to high-level attributes including hairstyle and expression; and the Detail Expert (DINOv2 [oquab2023dinov2]) highlights fine-grained textures and structural edges. These complementary roles indicate that robust identity preservation under large pose variations requires dynamic fusion of all three signals rather than reliance on a single static feature.
Table 1: Comparison with existing video datasets. Our LaFID-180K dataset uniquely provides identity-consistent and clustered human videos with skeleton and face pose annotations, supporting high-fidelity identity preservation in video generation.
Dataset name Domain # Videos Annotations Resolution ID-Consistent? ID-Clustered?
Koala-36M [wang2025koala] Open 36M - 720P
OpenHumanVid [li2025openhumanvid] Human 13.2M Skeleton pose 720P
LaFID-180K(Ours) Human 180K Skeleton pose, face pose 720P, 1080P
Collaborative Face Experts Fusion. Our goal is to fuse hierarchical face features adaptive to distinct DiT blocks, thereby achieving complementary synergies that enhance identity persistence, semantic coherence, and pixel-level fidelity. To achieve this, we first employ a face preprocessor comprising a face detector [deng2020retinaface] and a face parsing model [geitgey2019face] to obtain the cropped and aligned face image from the input image, removing background to focus on the face region. These preprocessed faces are then processed by identity, semantic, and detail encoders to generate identity-sensitive features, pixel-level detail features, and high-level semantic features, denoted as $\mathbf{f}_{\text{id}},\mathbf{f}_{\text{sem}},\mathbf{f}_{\text{det}}$, respectively.
The proposed CoFE module comprises two core components: i) Three globally shared projection layers, i.e., $P_{\text{id}}$, $P_{\text{sem}}$, $P_{\text{det}}$, which project $\mathbf{f}_{\text{id}}$, $\mathbf{f}_{\text{sem}}$, $\mathbf{f}_{\text{det}}$ into a unified feature space compatible with pre-trained DiT blocks, ensuring cross-modality alignment; ii) Per-DiT-block expert layers, each equipped with three dedicated experts, i.e., $\mathcal{E}_{\text{id}}$, $\mathcal{E}_{\text{sem}}$, $\mathcal{E}_{\text{det}}$, and a lightweight gating mechanism that dynamically fuses the projected features. Each expert layer is tailored to the semantic depth of its DiT block, using adaptive gating (via a learnable MLP) to prioritize experts based on the block’s role in denoising processing. This design ensures global feature consistency through shared projection layers and local adaptivity through block-specific expert weighting, enabling end-to-end optimization of cross-dimensional facial representations. Formally, each expert refines its input feature for the $i$-th DiT block as:
$\mathbf{e}^{i}_{\text{attr}}=\mathcal{E}^{i}_{\text{attr}}(P_{\text{attr}}(\mathbf{f}_{\text{attr}})),\text{attr}\in\{\text{id},\text{sem},\text{det}\}$
Then a learnable gate mechanism computes adaptive weights for these experts based on the concatenated input features $\mathbf{e}^{i}_{\text{c}}=[\mathbf{e}_{\text{id}}^{i},\mathbf{e}_{\text{sem}}^{i},\mathbf{e}_{\text{det}}^{i}]$:
$\mathbf{w}^{i}=\text{Softmax}\left(\mathcal{G}(\mathbf{e}^{i}_{\text{c}})\right),$
where $\mathcal{G}$ is a linear projection mapping $\mathbf{e}_{\text{c}}$ to weight vector $\mathbf{w}^{i}=[w^{i}_{\text{id}},w^{i}_{\text{sem}},w^{i}_{\text{det}}]$ with elements summing to 1. The final fused face feature, i.e., $\mathbf{f}^{i}_{\text{fused}}$, can be obtained via:
$\mathbf{f}^{i}_{\text{fused}}=\mathbf{w}^{i}\cdot\mathbf{e}^{i}_{c}$
DiT Feature Injection To integrate the fused face features with the DiT architecture, the output of the CoFE module, $\mathbf{f}^{i}_{\text{fused}}$, is injected into the cross-attention layers of the $i$-th DiT block, enabling video latents to learn face-specific features throughout the denoising process. This injection introduces a dedicated facial cross-attention stream that operates in parallel with DiT’s existing context and image cross-attention streams. Each stream contributes key and value vectors: $\mathbf{k}^{i}_{\text{ctx}},\mathbf{v}^{i}_{\text{ctx}}$ are derived from context, $\mathbf{k}^{i}_{\text{img}},\mathbf{v}^{i}_{\text{img}}$ are extracted from the first frame to preserve initial visual consistency, and $\mathbf{k}^{i}_{\text{fused}},\mathbf{v}^{i}_{\text{fused}}$ are projected from $\mathbf{f}^{i}_{\text{fused}}$ via learnable matrices to emphasize refined face features. Notably, the gate mechanism in CoFE dynamically adjusts the influence of the facial stream, adapting to various demands of different DiT blocks to achieve flexible face feature injection. The aggregated cross-attention output updates video latents as:
$\mathbf{o}^{i}$ $=\text{attn}^{i}(\mathbf{q}^{i},\mathbf{k}^{i}_{\text{ctx}},\mathbf{v}^{i}_{\text{ctx}})+\text{attn}^{i}(\mathbf{q}^{i},\mathbf{k}^{i}_{\text{img}},\mathbf{v}^{i}_{\text{img}})$
$\quad+\text{attn}^{i}(\mathbf{q}^{i},\mathbf{k}^{i}_{\text{fused}},\mathbf{v}^{i}_{\text{fused}}),$
where $\text{attn}(\cdot)$ denotes the attention function.
4 LaFID Benchmark
(a) Caption Word Count
(b) Caption Word Cloud
(c) Video Duration
(d) Video Resolution
Figure 4: The statistical analysis of the LaFID-180K dataset.
4.1 Overview
[…]
A key design goal is to ensure both consistent identity representation and rich dynamic pose variation. To this end, our processing pipeline employs objective facial quality filtering that naturally yields videos with high aesthetic appeal, strong imaging quality, and stable subject identity(Fig. 5). We quantify pose variation by estimating 3D face angles (pitch, yaw, roll) per frame and computing the maximum angular range within each clip. As Fig. 5(a) demonstrates, LaFID exhibits significantly broader pose coverage than existing datasets, especially in extreme yaw and pitch, making it uniquely suited for identity preservation under challenging face motions.
Existing benchmarks such as VIP-Test [vip-200k] focus on full-body motion but lack sufficient variation in large face poses, limiting their ability to evaluate identity preservation under challenging pose conditions. To address this gap, we introduce LaFID-Bench, a standardized evaluation protocol derived from LaFID-180K, comprising fifty held-out identities (disjoint from training) with four diverse prompts per identity, all featuring large face poses that span extreme yaw and pitch ranges. For generalization assessment, we additionally construct a VIP-Test subset using SDXL-generated [podell2023sdxl] reference images (fifty identities, four prompts each). Together, these two test sets enable the first holistic evaluation of image-to-video models across both standard full-body animation and the challenging regime of identity preservation under large face poses.
(a) Face Pose Variations
(b) Subject Consistency
(c) Aesthetic Score
(d) Imaging Quality
(e) Time Flickering
Figure 5: Comparison of video quality metrics between LaFID-180K and OpenHumanVid. LaFID-180K exhibits better quality.
Figure 6: The overview of the data processing pipeline for constructing dataset with large face poses.
4.2 Data Processing Pipeline
Our overall data processing pipeline is illustrated in Fig. 6. To collect videos with substantial face pose variations, we apply a multi-stage procedure. The first step, Face Constraints Filtering, retains clips exhibiting large face pose variations (e.g., extreme pitch, yaw or roll rotations). The second step, Identity Consistency Analysis, ensures facial identity remains coherent during pronounced face motion by discarding clips with inconsistent identity across frames. The third step, Speech Disambiguation, aligns lip motion or speech activity with caption by annotating whether the on-screen subject is speaking. Together, these steps ensure high-quality data for downstream identity-preserving tasks.
Table 2: Comparison with SOTA methods on LaFID-Bench benchmark dataset. The best results are highlighted in Bold ($\uparrow$$\downarrow$ represents the higher/lower the better). CoFE (Ours) has achieved the best performance across most metrics, demonstrating its superiority.
Method Identity Preservation Text Relevance Video Quality
FaceSim-Arc↑ FaceSim-Cur↑ CLIP-Score↑ FID↓ AQ↑ IQ↑ TF↑ MS↑
CogVideoX-5B [yang2024cogvideox] 0.629 0.636 29.331 104.214 0.499 0.598 0.976 0.986
LTX-2B [hacohen2024ltx] 0.438 0.440 28.800 151.773 0.513 0.581 0.980 0.992
ConsisID [yuan2025consisid] 0.567 0.573 31.184 123.967 0.560 0.670 0.979 0.990
Wan2.1-I2V-1.3B [wan2025wan] 0.654 0.656 29.200 72.322 0.553 0.660 0.989 0.993
Wan2.1 w/ CoFE 0.770 0.775 29.125 63.799 0.570 0.673 0.990 0.994
Wan2.2-TI2V-5B [wan2025wan] 0.776 0.786 28.761 76.728 0.588 0.666 0.995 0.996
Wan2.2 w/ CoFE 0.814 0.827 28.729 70.727 0.591 0.669 0.996 0.996
Table 3: Comparison with SOTA methods on the VIP-Test benchmark. Our CoFE markedly improves these metrics, achieving the best identity consistency while maintaining strong text and video quality.
Method Identity Preservation Text Relevance Video Quality
FaceSim-Arc↑ FaceSim-Cur↑ CLIP-Score↑ FID↓ AQ↑ IQ↑ TF↑ MS↑
CogVideoX-5B [yang2024cogvideox] 0.462 0.474 28.947 163.001 0.512 0.585 0.926 0.948
LTX-2B [hacohen2024ltx] 0.301 0.323 28.403 203.516 0.510 0.586 0.944 0.975
ConsisID [yuan2025consisid] 0.269 0.289 29.601 206.924 0.566 0.638 0.946 0.976
Wan2.1-I2V-1.3B [wan2025wan] 0.471 0.480 29.825 129.116 0.554 0.659 0.940 0.967
Wan2.1 w/ CoFE 0.502 0.510 30.167 126.676 0.558 0.679 0.947 0.971
Wan2.2-TI2V-5B [wan2025wan] 0.541 0.552 29.819 125.893 0.574 0.667 0.934 0.964
Wan2.2 w/ CoFE 0.605 0.615 29.834 113.850 0.583 0.668 0.948 0.979
Face Constraints Filtering. To support identity preservation under diverse face poses, especially large face motions, we curate high-quality samples for LaFID via multi-dimensional face constraints filtering, enforcing strict criteria on face continuity, visibility, and pose diversity. For efficiency, we sample one frame every three and detect face bounding boxes and poses using ArcFace [deng2020retinaface]. The core constraints are: i) Face count. Focused on single person generation, we filter videos where any selected frame contains no faces or multiple faces. This ensures a single, continuous face throughout the video clip, foundational for consistent identity learning. ii) Face proportion. Video clips with low face bounding box proportion (i.e., less than 10%) are filtered out to emphasize the preservation of facial details. iii) Face pose diversity. To ensure preservation of identity across all face poses, we retain videos with diverse poses. We analyze the face poses of selected frames for each video clip, selecting videos with large face pose variations.
Identity Consistency Analysis Identity consistency is essential for human-centric videos, reflecting the stability of facial representations across frames. To ensure high-quality data curation, we analyze identity coherence using ArcFace features [deng2019arcface]. i) Identity Similarity. We compute pairwise cosine similarities between frame embeddings $\{\mathbf{d}_{i}\}_{i=1}^{N}$:
$\text{sim}(\mathbf{d}_{i},\mathbf{d}_{j})=\frac{\mathbf{d}_{i}\cdot\mathbf{d}_{j}}{\|\mathbf{d}_{i}\|\|\mathbf{d}_{j}\|}\quad(i\neq j),$ (1)
and use the average off-diagonal similarity $\overline{s}$ as a measure of intra-clip identity coherence. Clips with $\overline{s}\leq 0.6$ are discarded. ii) Identity Clustering. To ensure disjoint identity splits for fair train/test evaluation, we cluster videos by identity using FAISS [douze2024faiss]. An IVF-PQ index [xiao2022distill] enables efficient similarity search in the ArcFace embedding space via product quantization. A two-stage hierarchical strategy is applied: a high threshold first forms tight identity clusters from strongly correlated samples; a lower threshold then assigns ambiguous samples to recover weakly correlated but same-identity instances.
Speech Disambiguation. LaFID is mainly curated from OpenHumanVid [li2025openhumanvid], where most videos show people speaking with continuous lip movements but captions rarely indicate actual speech or lip motion. Without explicit speech semantics, identity-preserving generators often wrongly link lip motion to text, producing videos with incessant lip movement even when no speech motion is contained in the input caption. To align lip motion with caption, we disentangle speaking status in two steps: whisper [radford2023robust] detects audible speech and its alignment with the primary subject’s lip movements; if the caption lacks explicit description of speaking, an LLM [bai2023qwen, touvron2023llama] appends a context-aware sentence (e.g., “The man is talking.”) to signal active speech. This ensures genuine speaking behavior is reflected in text-video alignment, enabling controllable and identity-consistent generation.
5 Experiments
Figure 7: Qualitative comparison of identity preservation and pose handling. For each sequence, we report face similarity (sim) to the reference (first-frame) identity and face pose deviation ($\Delta$, in degrees), defined as the angular distance between the current frame’s and the first frame’s 3D face orientations. CoFE achieves superior identity preservation and visual quality over existing methods.
5.1 Implementation Details
5.1.1 Training strategies
We integrate the proposed CoFE into the Wan2.1-I2V-1.3B and Wan2.2-TI2V-5B models [wan2025wan]. To balance performance and efficiency, we inject CoFE into each even-numbered DiT block, in line with VACE [jiang2025vace]. This design enhances model capabilities while incurring a modest parameter increase of merely 18.5%. We employ aspect ratio bucketing for multi-resolution and multi-frame training, with a minimum video resolution of 480p and a frame count ranging from 49 to 129. To address the “copy-paste” shortcut, we incorporate face features from non-training frames into the input for video clips with additional frames. We use the Adam optimizer with a learning rate of 1e-5 while keeping other modules frozen, and train on the LaFID-180K dataset for 4,000 iterations with a batch size of 256.
5.1.2 Evaluation Protocols
We evaluate our method on LaFID-Bench and VIP-Test [vip-200k], adopting the evaluation protocol of ConsisID [yuan2025consisid] for metric selection. Specifically, we assess identity preservation using FaceSim-Arc [deng2019arcface] and FaceSim-Cur [huang2020curricularface], text relevance via CLIP-Score [hessel2021clipscore], facial fidelity through FID, and overall video quality with VBench [huang2024vbench]. The latter includes four key sub-metrics: Aesthetic Quality (AQ), Imaging Quality (IQ), Temporal Flickering (TF), and Motion Smoothness (MS).
5.2 Benchmark Results
We compare our method with open-source state-of-the-art approaches [yang2024cogvideox, hacohen2024ltx, yuan2025consisid, wan2025wan] on LaFID-Bench and VIP-Test [vip-200k] to demonstrate the superiority of our CoFE module in preserving identity under large face poses.
5.2.1 Quantitative Evaluation
As shown in Tables 2 and 3, our CoFE delivers substantial improvements on both Wan2.1 and Wan2.2, with larger gains for Wan2.1 on LaFID-Bench and for Wan2.2 on VIP-Test. This consistent enhancement across model generations and generation scenarios reveals a key insight: identity preservation relies not on global video modeling alone but on the explicit injection of pose-robust, disentangled facial representations. Notably, our approach remains effective even in full-body settings where faces are small or temporally unstable, demonstrating that high-fidelity identity control can be achieved through localized facial priors decoupled from full-body motion synthesis.
On general video quality metrics (AQ, IQ, TF, and MS), CoFE outperforms all baselines, including Wan2.2, yielding higher visual appeal, fewer artifacts, and smoother motion. Existing methods exhibit distinct limitations: LTX-2B suffers from identity drift; ConsisID achieves high CLIP-Score but often generates blurry or distorted faces due to its narrow focus on facial alignment; and CogVideoX-5B delivers only moderate performance despite its large scale. Even the improved Wan2.2 falls short of CoFE in both identity fidelity and overall quality. These results demonstrate that explicitly modeling disentangled facial attributes such as identity, semantics, and detail as structured priors enables more robust and coherent video generation under complex dynamics.
Table 4: Comparison of face feature injection methods on LaFID-Bench. CoFE achieves the best identity preservation while maintaining strong text relevance and fidelity.
Method FaceSim-Arc↑ FaceSim-Cur↑ CLIP-Score↑ FID↓
Wan-I2V [wan2025wan] 0.654 0.656 29.200 72.322
w/ Identity Expert 0.661 0.665 29.402 80.305
w/ Semantic Expert 0.681 0.684 29.181 73.702
w/ Detail Expert 0.672 0.675 29.079 75.467
w/ Concat 0.732 0.735 29.143 63.418
w/ CoFE 0.770 0.775 29.125 63.799
Figure 8: Qualitative comparison between CoFE module and other face feature injection methods. Our proposed CoFE demonstrates superior identity preservation under large face angle variations in video sequences.
5.2.2 Qualitative Evaluation
Fig. 7 presents qualitative comparison between our method and other SOTA methods. In scenarios with significant face movements, our model exhibits more stable generation and superior identity preservation: multi-source feature injection enhances facial details, ensuring consistent identity throughout the video clip. Specifically, ConsisID poorly renders front-facing details when given only a profile as input. CogVideoX and Wan2.2 fail to preserve identity during large face movements, causing obvious distortions and identity shifting. LTX-Video generates unstable, often blurry frames with lost identity details. Supplementary videos further validate our approach’s robustness.
5.3 Ablation Analysis
Feature Fusion Strategy Comparison. To validate the advantages of our adaptive face feature fusion, we compare CoFE with two common baselines: single-expert injection using the identity, semantic, or detail expert alone, and naive concatenation of all three experts. We evaluate these strategies on LaFID-Bench using Wan2.1-I2V-1.3B as the base model. As shown in Table 4 and Fig. 8, single-expert variants improve over the baseline but exhibit distinct weaknesses: the identity expert lacks fine details such as eye shape, the semantic expert shows inconsistent skin tones, and the detail expert yields blurry contours under large poses. Naive concatenation aggregates cues but treats them uniformly, introducing redundancy and misalignment that harm identity consistency. In contrast, CoFE achieves the highest FaceSim-Arc and FaceSim-Cur scores and the lowest FID, confirming that adaptive fusion of complementary facial representations is key to robust identity preservation.
Table 5: Ablation study on training datasets. Improvements from Pose Filtering and Speech Disambiguation (SD) are evident: LaFID w/o SD outperforms OpenHumanVid, and full LaFID further boosts identity fidelity and semantic alignment.
Dataset FaceSim-Arc↑ FaceSim-Cur↑ CLIP-Score↑ FID↓
OpenHumanVid 0.801 0.813 28.659 75.578
LaFID w/o SD 0.807 0.816 28.634 73.128
LaFID 0.814 0.827 28.729 70.727
Ablation of Training Data Processing. In our data processing pipeline, we introduce Face Constraints Filtering and Speech Disambiguation to enhance identity stability and semantic alignment. As shown in Table 5, models trained on our filtered data consistently outperform those using the raw OpenHumanVid subset, validating the effectiveness of our proposed data processing pipeline. Furthermore, incorporating Speech Disambiguation yields additional gains across all metrics, demonstrating that explicitly aligning captions with speaking status helps mitigate spurious lip-motion cues and improves identity preservation. More qualitative results are provided in our supplementary material.
6 Conclusion
In this paper, we address facial identity preservation under large pose variations in image-to-video generation through two key contributions. First, we introduce the Collaborative Face Experts Fusion (CoFE) module, which dynamically fuses identity-sensitive, semantic, and detail-oriented face features through a layer-adaptive cross-attention gating mechanism. Second, we curate LaFID-180K through a three-stage pipeline: facial constraints filtering for large-pose diversity, identity consistency analysis for cross-frame identity stability, and speech disambiguation to align captions with speaking status, effectively reducing lip motion and text ambiguity. To support rigorous evaluation in the challenging regime of large face poses, we introduce LaFID-Bench, a new benchmark specifically designed for this scenario. Experiments show that CoFE achieves state-of-the-art performance on both the general-purpose VIP-Test [vip-200k] and our specialized LaFID-Bench, demonstrating consistent improvements in identity preservation and overall video quality across diverse generation settings.